# Evaluating statistical significance for massive black hole binary mergers with space-based gravitational wave detectors

### v1

@@ Abstract
Important scientific discoveries should be backed by high statistical significance. In the 2030s, multiple space-based gravitational wave detectors are expected to operate. While many works aim to achieve quick and reliable detection and parameter estimation of millihertz gravitational wave sources, dedicated studies are lacking to assess the significance of space-based detectors. In this work, we propose a framework to assess the statistical significance of massive black hole binaries (MBHBs) detections with space-based gravitational wave detectors. We apply this algorithm to simulated data with Gaussian stationary noise and the complex LDC-2a dataset to measure the false alarm rate and significance of MBHB signals. We also analyze factors affecting the significance of MBHBs and design a method to mitigate multi-source confusion interference. In Gaussian noise conditions, MBHBs with a signal-to-noise ratio of about 7 can achieve $3\sigma$ significance, and those with a signal-to-noise ratio of about 8 achieve $4\sigma$. Our analysis demonstrates that all MBHB signals in the LDC-2a dataset have a significance exceeding $4.62\sigma$.
@@ I Introduction
This work aims to develop a significance evaluation algorithm suitable for space-based GW detectors, specifically for the detection of MBHB mergers. In this study, we will select suitable time-shift intervals for space-based GW detectors and construct a significance evaluation pipeline based on the iterative Nelder-Mead algorithm.
@@
This paper is organized as follows: Section II gives an overview of the theoretical background of significance and its evaluation. In Section III, we analyze both the simulated Gaussian noise data and the LDC-2a Sangria dataset separately. Finally, Section IV summarizes the entire work. Appendix A provides a brief introduction to future space-based GW detectors. Appendix B presents the MBHB waveforms used in this work.
@@ II.1 Statistical significance
Achieving such high precision FAP is challenging. The time-shift method effectively addresses this by offsetting the two detectors’ timestamps to decorrelate the data and construct long-duration equivalent background data. For ground-based detector networks, using this method with several days of observational data from three detectors allows for constructing background data spanning longer than the universe’s age [41]. Then, FAR and FAP can be easily calculated by searching for coincident triggers in the equivalent background data:
${\rm FAR}$ $=\frac{N_{\rm FA}}{T_{\rm tot}},$ (1)
${\rm FAP}$ $=\frac{T_{\rm obs}\times N_{\rm FA}}{T_{\rm tot}},$ (2)
where $N_{\rm FA}$ indicates the number of coincident triggers in the equivalent background data, $T_{\rm tot}$ and $T_{\rm obs}$ are the total time of equivalent background data and the actual observation time.
@@ II.2 Assessment of significance
For space-based GW detectors, we can adapt the time-shift concept with targeted optimizations, as depicted in 1. Unlike ground-based detectors that require multiple detectors for time-shifting, a triangular-shaped space-based GW detector includes three interferometers. These data streams can be rearranged into three noise-uncorrelated channels: $A$, $E$, and $T$ (see Equation 5). Due to the $T$ channel’s low sensitivity, which barely affects the time-shift algorithm’s results, we focus on $A$ and $E$. After whitening, we can construct equivalent background noise from the $A$ and $E$ channels. Considering the light-second scale of space-based detectors, as well as the timing errors, we’ve chosen a 1-minute interval for time-shifting. This choice is a compromise so that the two data channels are uncorrelated after the time-shift, meanwhile, one can construct large enough background noise.
@@
Time-shift algorithms in ground-based GW detectors contain two assumptions: inclusive background and exclusive background. The latter excludes all triggers with candidates that have a FAR below a specified threshold, thus offering a more conservative description of the noise background distribution [42, 43]. For space-based GW detectors, MBHB signals, due to higher SNRs and longer durations, have a more significant impact on the background distribution. Consequently, this study focuses solely on analyzing the exclusive background from which MBHB signals have been removed.
When identifying triggers in time-shifted data, we used the Nelder-Mead algorithm instead of the traditional template bank method. The latter is computationally impractical for higher-dimensional, longer-duration signals from space-based GW detectors [44, 45]. In contrast, the Nelder-Mead algorithm, a derivative-free multidimensional optimization method [46, 47], exceeds most search algorithms in performance. It needs minimal function evaluations per iteration, usually one or two, making it ideal for resource-heavy function evaluations [48]. Its efficiency and ability to handle non-smooth or discontinuous functions without derivative information make it a suitable choice for fast optimization in practical applications. Though sometimes trapped in local rather than global maxima, iterative algorithms can still identify all posterior probability peaks in data.
In the selection of the detection statistic, we chose an SNR form based on the log-likelihood ratio.
$\rho$ =$=\left\{\begin{matrix}\sqrt{2\times\mathcal{\rm ln\Lambda}},&\mathcal{\rm\Lambda}>0\\ 0,&\mathcal{\rm\Lambda}\leq 0\end{matrix}\right.$ (3)
$\mathcal{\rm ln\Lambda}$ $=\langle d\mid h\rangle-\frac{1}{2}\langle h\mid h\rangle,$ (4)
where, $d$ is the data, and $h$ is the signal template. This form of SNR is better at suppressing triggers caused by noise compared to the optimal SNR in data analysis [49]. For the construction of waveform $h$, see Appendix B.
@@ III.1 Gaussian Noise
Figure 2: The left panel shows the theoretical noise $A/E$ channel PSD of TianQin (red) and LISA (blue). The right panel illustrates the SNR distribution of false alarms. The left-hand axis of the right panel shows the cumulative count rate (FAR) at different SNRs. The right-hand axis is the significance corresponding to different FARs. The case for LISA with the foreground is marked with the blue dashed line.
@@
By shifting 1-month observational data at 1-minute intervals, we moved the data 43,200 times to create equivalent background data spanning 3,600 years. For each piece of data, by applying the Nelder-Mead algorithm, we can identify the trigger with the highest SNR within the 1-month period. All triggers with SNR greater than 3 are accumulated and their distribution shown in Fig. 2(b). Since no authentic signals are expected in the time-shifted background, all the triggers can be treated as false alarms. By setting an SNR threshold, one can obtain the FAR as the cumulative count rate of the false alarms which are associated with SNR larger than the threshold. Based on Table 1, we calculated the relationship between FAR and significance for a 1-month observation. The lowest FAR of $1/3600~{}{\rm yr}^{-1}$ corresponds to a significance level of $4.07\sigma$. In Fig. 2(b), we mark the cumulative count rate and significance for MBHB signals with different SNRs.
Based on the SNR distribution of false alarms from Gaussian noise, Table 2 presents the SNR required for TianQin and LISA (both the instrumental noise itself and that includes foreground noise) to reach $3\sigma$ and $4\sigma$. TianQin has a lower FAP for MBHB detection. For LISA, the FAP slightly increases when foreground noise is included.
@@
To explore the causes of false alarms and detector-specific FAP differences, we collected false alarms with $\rho\geq 3$. These false alarms are mostly uniformly and randomly distributed in the parameter space but show patterns in chirp mass and luminosity distance. As shown in Fig. 3, most false alarms involve higher or lower masses and greater distances. Luminosity distance affects GW amplitude, while chirp mass determines GW frequency. Since the data are Gaussian noise based on detector PSD, in frequency bands with low PSD, noise and random fluctuations are minimal, leading to weak false-alarm triggers and a low FAP. Consequently, TianQin and LISA have extremely low FAPs in their sensitive bands (corresponding to MBHBs with chirp masses about $\sim 10^{5}~{}M_{\odot}$ and $\sim 10^{6}~{}M_{\odot}$). In contrast, in bands with higher noise PSD, their FAPs are higher. Meanwhile, TianQin’s superior sensitivity in higher frequencies gives it a low FAP at the low-mass end. As for the high-mass end, the $10^{-4}$ Hz truncation of TianQin hinders false alarms from accumulating enough SNR, also keeping the FAP low. For LISA, when considering the foreground, many false alarms appear with a chirp mass of about $10^{6}~{}M_{\odot}$. This corresponds to the noise PSD increase in LISA due to foreground noise at around 1 mHz.
@@
Figure 3: The histograms of the distribution of false alarms with $\rho\geq 3$ in terms of chirp mass and luminosity distance for different detectors.
@@ III.2 LDC-2a Sangria Dataset
To reduce contamination from detected signals, we obtained the MBHB parameters in the Sangria dataset via low-latency detection from Cornish [34]. Figure 4 shows the 15 detected signals, with the x-axis representing the SNR under the Smooth PSD, the y-axis showing the relative deviation of SNR, $\Delta\rho/\rho({\rm Smooth~{}PSD})$, and the color indicating the chirp mass. Only one signal has an SNR deviation exceeding 10% in the two PSD conditions. After subtracting these signals from the original dataset, we analyzed the residuals. By conducting $525,600$ time-shifts, we generated hundreds of thousands of years of equivalent background data. The lowest FAR of $1/525,600~{}{\rm yr}^{-1}$ corresponds to a significance level of $4.62\sigma$.
Figure 4: The SNR of 15 detected signals under the Smooth PSD condition and their deviations in two PSD cases. The chirp mass of the signals is indicated by color.
@@
Figure 6: The histograms of the distribution of false alarms, similar with Fig. 3, but for Sangria dataset and the SNR threshold is 5.
The distribution of false alarms provides interesting insights. As shown in Fig. 6, which illustrates the false alarm distribution in both scenarios, the mass-distance distribution of false alarms under the Spectrum PSD closely resembles that of LISA Gaussian noise with foreground in Fig. 3. This shows that the Spectrum PSD effectively suppresses the impact of high SNR GBs on MBHB significance. In contrast, the Smooth PSD results in significantly more false alarms with chirp masses of $10^{5}\sim 10^{6}~{}M_{\odot}$.
Furthermore, Fig. 6 indicates that false alarms are not expected to be linked with nearby events ($\lesssim 50{\rm Gpc}$) in the mass range of $10^{5}\sim 10^{7}~{}M_{\odot}$. Therefore, MBHB candidates detected in this range are expected to be associated with higher significance, and it is possible to use this information to further boost their significance, so that even relatively weak signals could be unambiguously identified.
@@
Table 3: The SNR required for MBHBs to reach a specified significance level, similar with Table 2, but under LDC-2a Sangria dataset with two kinds of PSDs.
@@ IV Summary
Evaluating significance is essential to establish faith in future milli-Hertz GW detections. Yet, studies on implementing significance estimation algorithms are still inadequate. In this work, we develop a fast, robust, and accurate significance evaluation algorithm for MBHB mergers with space-based GW detectors. We leverage the time-shift method and the iterative Nelder-Mead algorithm, combined with the $A$, $E$, and $T$ channel characteristics of space-based detectors, to achieve significance evaluation for MBHBs under both ideal conditions and multi-source confusion scenarios.
@@
We applied our significance evaluation pipeline to simulated Gaussian noise data and the LDC-2a dataset, and analyzed how factors like foreground noise and GB signals affect the significance of MBHBs. For the LDC-2a dataset with multi-source confusion, we found that treating high SNR GBs as spikes in the PSD can greatly mitigate their contamination in false alarms, so that the significance of actual MBHBs can be greatly enhanced. Our analysis of one month of Gaussian and stationary data takes several thousand core hours, whereas analyzing a year of LDC-2a data requires tens of thousands of core hours. However, given the code’s high parallelizability, the actual wall-clock time remains manageable in a multi-core parallel computing environment.
In simulated Gaussian noise data, the significance of MBHB systems with SNR about $7\left(8\right)$ on TianQin or LISA can reach $3\sigma\left(4\sigma\right)$. False alarms are mostly uniformly distributed across parameters but exhibit patterns in chirp mass and luminosity distance. Fewer false alarms occur in detector-sensitive frequency bands, and they tend to cluster at the greatest distances. In addition, we have taken into account the potential impact of foreground noise on LISA. Furthermore, we considered the potential impact of foreground noise on LISA. Specifically, around 1 mHz of foreground noise, LISA’s simulated Gaussian noise data generated numerous false alarms with chirp masses of $10^{6}~{}M_{\odot}$.
In the LDC-2a dataset, we analyzed the exclusive background using two different PSDs. Among the 15 detected MBHB sources, the maximum deviation caused by the two PSDs was only approximately 10%. However, the Spectrum PSD can significantly reduce the signal’s FAR by several orders of magnitude, lowering the required SNR for a $4\sigma$ detection from around 35 to about 10. The Spectrum PSD effectively suppressed false alarms with chirp masses of $10^{5}\sim 10^{6}~{}M_{\odot}$, which aligns with the main frequency band of the GB signals. Consequently, under the Spectrum PSD, the relationship between significance and SNR, as well as the distribution of false alarms, resembles that of Gaussian noise data with foreground noise. In contrast, the Smooth PSD led to numerous GB signals being misidentified as MBHBs, significantly increasing the FAR. Nevertheless, due to the relatively high SNR of MBHBs, in both cases, no false alarm SNR exceeded that of the detected signals. Thus, all 15 detected signals achieved a significance level exceeding $4.62\sigma$.
@@
Our current algorithm struggles to achieve $5\sigma$ significance, unless given longer observation times or smaller time-shift intervals. The analysis is also based on the assumption that noise PSD is given or at least uncorrelated. We plan to address these limitations in future work [51]. Enhancement is feasible, and we hereby propose several potential approaches to resolve this issue. First, capitalizing on the unique properties of MBHB signals, such as the SNR accumulation patterns [52] and the mass-distance relationship, can further refine the filtering process. This approach eliminates potential false alarms and boosts the significance of candidates matching these features. Second, future multi-detector network will exponentially increase time-shift combinations via more observational channels, enabling a longer equivalent background data construction and lower FAP limit. Third, using a Poisson estimation method, we can extrapolate and estimate the significance of higher SNR signals based on the existing significance-SNR relationship.

### v2

# arXiv:2505.09446v2[gr-qc] 01 Mar 2026

@@ Abstract
Important scientific discoveries should be backed by high statistical significance. In the 2030s, multiple space-based gravitational wave detectors are expected to operate. While many works aim to achieve quick and reliable detection and parameter estimation of millihertz gravitational wave sources, dedicated studies are lacking to assess the significance of space-based detectors. In this work, we propose a framework to assess the statistical significance of massive black hole binary (MBHB) detections with space-based gravitational wave detectors. We apply this algorithm to simulated data with Gaussian stationary noise and the complex LDC-2a dataset to measure the false alarm rate and significance of MBHB signals. We also analyze factors affecting the significance of MBHBs and design a method to mitigate multi-source confusion interference. In Gaussian noise conditions, MBHBs with a signal-to-noise ratio of about 7 can achieve $3\sigma$ significance, and those with a signal-to-noise ratio of about 8 achieve $4\sigma$. Our analysis demonstrates that all MBHB signals in the LDC-2a dataset have a significance exceeding $4.62\sigma$.
@@ I Introduction
This work aims to develop a significance evaluation algorithm suitable for space-based GW detectors, specifically for the detection of MBHB mergers. The estimate of significance is contingent on the search method, and we aim to push the limit with original data, with minimal assumptions on the statistical properties of the noise. In this study, we will select suitable time-shift intervals for space-based GW detectors and construct a significance evaluation pipeline based on the iterative Nelder-Mead algorithm.
@@
This paper is organized as follows: Section II gives an overview of the theoretical background of significance and its evaluation. In Section III, we analyze both the simulated Gaussian noise data and the LDC-2a Sangria dataset separately. Finally, Section IV summarizes the entire work. Appendix A provides a brief introduction to future space-based GW detectors.
@@ II.1 Statistical significance
Achieving such high precision FAP is challenging. The time-shift method effectively addresses this by offsetting the two detectors’ timestamps to decorrelate the data and construct long-duration equivalent background data. For ground-based detector networks, using this method with several days of observational data from three detectors allows for constructing background data spanning longer than the universe’s age [20]. Then, FAR and FAP can be easily calculated by searching for coincident triggers in the equivalent background data: {align} FAR = NFATtot,
FAP = Tobs×NFATtot, where $N_{\rm FA}$ indicates the number of coincident triggers in the equivalent background data, $T_{\rm tot}$ and $T_{\rm obs}$ are the total time of equivalent background data and the actual observation time.
@@ II.2 Assessment of significance
For space-based GW detectors, we can adapt the time-shift concept with targeted optimizations, as depicted in 1. Unlike ground-based detectors that require multiple detectors for time-shifting, a triangular-shaped space-based GW detector includes three interferometers. These data streams can be rearranged into three noise-uncorrelated channels: A, E, and T [32, 53]. Due to the T channel’s low sensitivity, which barely affects the time-shift algorithm’s results, we focus on $A$ and $E$. After whitening, we can construct equivalent background noise from the A and E channels. Considering the light-second scale of space-based detectors, as well as the timing errors, we’ve chosen a 1-minute interval for time-shifting. This choice is a compromise so that the two data channels are uncorrelated after the time-shift, meanwhile, one can construct a large enough background noise. We remark that although for space-borne missions, glitches can be correlated and thus corrupt the sensitivity, they are less common and come from fewer sources than their ground-based counterparts [40, 52, 8], so their effect are less severe. Thanks to the steady improvement in the glitch-cleaning pipeline [45, 10], it can be expected that all major glitches can be identified and gated without harming channel independence. Of course, this assumption was adopted for simplicity, and weaker glitches might be much harder to identify and could complicate the analysis. Noise-monitoring channel can be used to flag the glitches [55], and, in the worst case, we can perform the time slide on channels from different detectors like TianQin and LISA.
@@
Time-shift algorithms in ground-based GW detectors contain two assumptions: inclusive background and exclusive background. The latter excludes all triggers with candidates that have a FAR below a specified threshold, thus offering a more conservative description of the noise background distribution [2, 16]. For space-based GW detectors, MBHB signals, due to higher SNRs and longer durations, have a more significant impact on the background distribution. To eliminate this contamination and safely adopt a shorter time-shift interval, we confine our analysis to the exclusive background, i.e., the data set from which every identified MBHB signal has been removed.
When identifying triggers in time-shifted data, we used the Nelder-Mead algorithm, a derivative-free multidimensional optimization method [41, 33], which exceeds most search algorithms in speed. It iteratively refines a simplex via four core operations, reflection, expansion, contraction, and shrinkage, to approach the optimal solution. Instead of exploring the whole parameter space with a template bank, we use the Nelder-Mead method to iteratively refine an N-dimensional simplex, which converges to a maximum of the likelihood. The method is quick since only a limited number of likelihood evaluations are required. Notice that it comes with a price that the method does not guarantee a global maximum and is dependent on the choice of the initial simplex [49]. In practice, a single GW search with 500 iterations consumes about a few core minutes on a core for one year of data. Its efficiency and ability to handle non-smooth or discontinuous functions without derivative information make it a suitable choice for fast optimization in practical applications. To address the Nelder-Mead algorithm’s sensitivity to initial values and susceptibility to local maxima, we optimize it via iterative loops, rendering it robust to initial value variations. Specifically, the standard algorithm is integrated into an iterative subtraction loop: post convergence, peaks with SNR exceeding the threshold are removed from the data; otherwise, the search restarts with random initial values. The process terminates upon multiple consecutive failures to detect new extrema, thus identifying all statistically significant peaks.
In the selection of the detection statistic, we chose an SNR form based on the log-likelihood ratio. {align} ρ= {2×lnΛ, lnΛ¿0
0, lnΛ≤0
lnΛ= ⟨d ∣h ⟩- 12 ⟨h ∣h ⟩, where, $d$ is the data, and $h$ is the signal template. This form of SNR is better at suppressing triggers caused by noise compared to the optimal SNR in data analysis [11]. In this work, we utilized the IMRPhenomD template [27, 31, 22] and time delay interferometry (TDI) response [53, 39, 34] to generate MBHB signals template. In this work, we restrict the chirp mass range to $10^{5}\sim 10^{8}~M_{\odot}$ and set an upper limit on the luminosity distance of 230 Gpc, which corresponds approximately to a redshift of 20. This choice of upper limit is motivated by two factors: first, this range is around the edge of the horizon distance of space-borne detectors; secondly, astronomical models do not support mergers of MBHBs at larger redshifts. Therefore, triggers from events with a larger redshift could only arise from false alarms, which usually should be associated with low SNR. We believe that the choice of this upper limit in distance should not significantly alter our conclusions. For the other parameters, we draw the starting points uniformly.
@@ III Result
In the ideal case, where the signals can be exactly matched and removed, and when the noise is Gaussian and stationary, the statistical distribution of the false alarms can be analytically obtained. In this section, we first test our algorithm in the Gaussian and stationary noise case to verify if the false alarm distribution of the Nelder-Mead search follows the Rayleigh distribution. Then we apply the algorithm to the LDC-2a data set that contains multiple types of overlapping signals to test the robustness.
@@ III.1 Gaussian Noise
Figure 2: The left panel shows the theoretical noise A/E channel PSD of TianQin (red), LISA (blue), and LISA with foreground (green). The right panel illustrates the SNR distribution of false alarms, which is compared with the survival function of the Rayleigh distribution (lighter areas). The blue dashed line denotes the results of LISA after the low-frequency hard cutoff. The left-hand axis of the right panel shows the cumulative count rate (FAR) at different SNRs. The right-hand axis is the significance corresponding to different FARs.
@@
By shifting 1-month observational data at 1-minute intervals, we moved the data 43,200 times to create equivalent background data spanning 3,600 years. For each piece of data, by applying the Nelder-Mead algorithm, we can identify the trigger with the SNR greater than 3. All triggers with SNR greater than 3 are accumulated and their distribution is shown in Fig. 2(b). Since no authentic signals are expected in the time-shifted background, all the triggers can be treated as false alarms. By setting an SNR threshold, one can obtain the FAR as the cumulative count rate of the false alarms which are associated with SNR larger than the threshold. In Fig. 2(b), we mark the cumulative count rate and significance for MBHB signals with different SNRs. Our results closely match the expected Rayleigh survival function, confirming the reliability of the proposed algorithm. The lowest FAR of $1/3600~{\rm yr}^{-1}$ corresponds to a significance level of $4.07\sigma$.
Based on the SNR distribution of false alarms from Gaussian noise, Table 2 presents the SNR required for TianQin and LISA (both the instrumental noise itself and that includes foreground noise) to reach $3\sigma$ and $4\sigma$. TianQin has a lower FAP for MBHB detection. After investigation, we conclude that such a difference stems from the difference in the low-frequency cutoff between the two missions. Compared with TianQin ($10^{-4}$Hz), LISA has a lower low-frequency cutoff for the frequency band ($10^{-5}$Hz). There are only a limited number of data points between the range of $10^{-5}-10^{-4}$Hz, and random fluctuations from Gaussian noise could fool the search algorithm into treating it as containing a low-SNR, high-mass MBHB merger. The occurrence rate for lower-mass false alarms would be much smaller since it corresponds to higher frequencies and thus more data points are involved, making it harder for pure noise to mock the signal. To verify this hypothesis, we applied a hard cutoff to LISA’s low-frequency data at $10^{-4}$ Hz. It was found that the low-SNR false alarm rate of LISA decreased by an order of magnitude, approaching the results of TianQin. This is clearly illustrated by the blue dashed line in Fig. 2(b). We can observe that after the hard cutoff, the false alarms at low SNRs of LISA are roughly comparable to those of TianQin, indicating that the excess of low-SNR false alarms of LISA primarily stems from Gaussian random fluctuations in the low-frequency data. Meanwhile, in the high-SNR end, the dashed line approaches the solid line, indicating that the low-frequency data has a negligible contribution.
@@ III.2 LDC-2a Sangria Dataset
To reduce contamination from detected signals, we obtained the MBHB parameters in the Sangria dataset via low-latency detection from Cornish [19]. SNR deviations ($\Delta\rho/\rho({\rm Smooth~PSD})$) of most signals under two PSD conditions are under 8%, with the maximum deviation being 10.6%. After subtracting these signals from the original dataset, we analyzed the residuals. By conducting $525,600$ time-shifts, we generated hundreds of thousands of years of equivalent background data. The lowest FAR of $1/525,600~{\rm yr}^{-1}$ corresponds to a significance level of $4.62\sigma$.
@@
Figure 4: The histograms of the distribution of false alarms with $\rho\geq$ 5 (left panels) or 8 (right panels) in terms of chirp mass and luminosity distance for different PSDs.
To explore the causes of false alarms, we collected false alarms with $\rho\geq$ 5 or 8. These false alarms are mostly uniformly and randomly distributed in the parameter space but show patterns in chirp mass and luminosity distance. Fig. 4 shows that the SNR of false alarms in the high-mass regime, primarily attributable to random fluctuations in the noise, is consistently below 8. This conclusion is consistent with the results presented earlier (Fig. 2(b)). Moreover, the Spectrum PSD effectively suppresses the impact of high SNR GBs on MBHB significance. In contrast, the Smooth PSD results in significantly more false alarms with chirp masses of $10^{5}\sim 10^{6}~M_{\odot}$.
Furthermore, Fig. 4 indicates that false alarms are not expected to be linked with nearby events ($\lesssim 50{\rm Gpc}$). Therefore, MBHB candidates detected in this range are expected to be associated with higher significance, and it is possible to use this information to further boost their significance, so that even relatively weak signals could be unambiguously identified.
@@
Table 3: The SNR required for MBHBs to reach a specified significance level, similar to Table 2, but under the LDC-2a Sangria dataset with two kinds of PSDs.
@@ IV Summary
Evaluating significance is essential to establish faith in future milli-Hertz GW detections, while the estimation is tightly bound to the search algorithm adopted. Yet, studies on implementing significance estimation algorithms are still inadequate. In this work, we develop a fast, robust, and accurate significance evaluation algorithm for MBHB mergers with space-based GW detectors. We leverage the time-shift method and the iterative Nelder-Mead algorithm, combined with the $A$, $E$, and $T$ channel characteristics of space-based detectors, to achieve significance evaluation for MBHBs under both ideal conditions and multi-source confusion scenarios.
@@
We applied our significance evaluation pipeline to simulated Gaussian noise data and the LDC-2a dataset, and analyzed how factors like foreground noise and GB signals affect the significance of MBHBs. For the LDC-2a dataset with multi-source confusion, we found that treating high SNR GBs as spikes in the PSD can greatly mitigate their contamination in false alarms, enhancing the significance of actual MBHBs. Our analysis of one month of Gaussian and stationary data takes several thousand core hours, whereas analyzing a year of LDC-2a data requires tens of thousands of core hours. However, given the code’s high parallelizability, the actual wall-clock time remains manageable in a multi-core parallel computing environment.
@@
It should be noted that no glitches have been incorporated into the data of this work. In future work, we will conduct further investigations into the impacts of glitches on the FAR and potential mitigation strategies. In addition, our current algorithm struggles to achieve $5\sigma$ significance unless given longer observation times or smaller time-shift intervals. Enhancement is feasible, and we hereby propose several potential approaches to resolve this issue. First, capitalizing on the unique properties of MBHB signals, such as the SNR accumulation patterns [18] and the mass-distance relationship, can further refine the filtering process. This approach eliminates potential false alarms and boosts the significance of candidates matching these features. Second, the future multi-detector network will exponentially increase time-shift combinations through more observational channels, enabling a longer equivalent background data construction and a lower FAP limit. Third, using a Poisson estimation method, we can extrapolate and estimate the significance of higher SNR signals based on the existing significance-SNR relationship. Moreover, once we have a clear understanding of the noise properties, we can use this knowledge to model and simulate the noise for significance evaluation.